OpenStudy (jjuden):

Which equation would best help solve the following problem? Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m^2 by increasing the width and length by the same amount. What will be the dimensions of the new garden?

OpenStudy (jjuden):

a 208 = (10)(13) b 208 = (10 + x)(13 + x) c 208 = (10 +13)x d 208 = (10 + x^2)(13)

OpenStudy (anonymous):

Try B